# F3D-Gaus: Feed-forward 3D-aware Generation on ImageNet with Cycle-Consistent Gaussian Splatting

### v1

@@ 3.3 Geometry-guided Texture Refinement
We found that although our cycle-consistent self-supervised strategy improves 3D representations in novel views, there are still artifacts near edges with significant viewpoint changes. To address this, we introduce a fixed video in-painting model [69] in the second training stage for artifact localization and additional refinement.
Artifacts localization using normal map. When rendering novel view images from $\hat{GS_{0}}$, we also render the alpha map and calculate the normal map from the depth [63]. We found that alpha and normal maps effectively help identify artifacts. Specifically, if the alpha value is low and the angle between the normal and viewing direction is small, it is likely an artifact. This can be formalized as follows:
@@
where $I_{\text{in}}^{m}$ represents the in-painted image for the $m$-th frame, $\hat{I^{m}}$ is the rendered image, and $M^{m}$ is the mask generated from Equ. 12. Fig. 4 (b) shows one view of the in-painted results. Then, the corresponding video in-painting loss is defined as:
@@ 4.1 Experimental Setup
Dataset. We conducted experiments on the ImageNet dataset [10], Dogs [27], SDIP Elephants [27], and LSUN Horses [61], following previous methods. ImageNet contains 1000 categories of images, while the other three datasets consist of single-class images. For a fair comparison, we utilize pseudo-ground-truth depth from LeReS [58], though our approach can be extended to other state-of-the-art depth estimation methods. We follow the 3DGP [38] settings for ImageNet, training on a filtered subset (420K images) and evaluating the full training set (1.2M images). Most of our experiments are conducted on ImageNet.
Metrics. For the generated images, we compute Fréchet Inception Distance (FID) [16] and Inception Score (IS) on the full ImageNet training set. For the predicted depth maps, we calculate the normalized L2 distance between the predicted depth and pseudo ground-truth depth. Additionally, we compute the Non-Flatness Score (NFS) [38], a no-reference depth quality metric that evaluates the continuity by analyzing the histogram distribution of the depth map.
Method Synthesis FID $\downarrow$ IS $\uparrow$
BigGAN [1] 2D 8.7 142.3
StyleGAN-XL [33] 2D 2.3 265.1
ADM [11] 2D 4.6 186.7
EG3D [3] 3D-aware 25.6 57.3
StyleNeRF [14] 3D-aware 56.5 21.8
3DPhoto [36] 3D-aware 116.6 9.5
EpiGRAF [37] 3D 58.2 20.4
3DGP [38] 3D 19.7 124.8
VQ3D [31] 3D 16.8 n/a
G3DR [30] 3D 13.1 151.7
Our F3D-Gaus 3D 1.6 308.6
Table 1: Fidelity comparison on ImageNet $256^{2}$. Our F3D-Gaus surpasses all listed state-of-the-art 3D generation methods in both FID and IS metrics, achieving results comparable to those of 2D generation methods.
@@ 4.2 Main Results
Quantitative results. Fig. 5, Tab. 1, Tab. 2 and Tab. 3 present the qualitative and quantitative results of our method on the ImageNet dataset. The metrics in Tab. 1 and Tab. 2 are computed on the full training set at a resolution of 256, where our method outperforms all baselines across all metrics. In Tab. 1, single-image reconstruction methods are grouped into 2D and 3D approaches. Generally, 2D methods achieve higher scores than 3D methods, with FID values typically below 10. Among 3D methods, our approach significantly outperforms the others, achieving an FID 1.6 and an IS 308.6, reaching a performance level comparable to 2D methods. This indicates that our generated images are closer to real images in distribution. Similarly, in Tab. 2, our method achieves the best results, demonstrating superior performance in rendering depth maps. Tab. 3 shows our method significantly outperforms the baseline in image fidelity at a resolution of 128.
Method Depth accuracy $\downarrow$ NFS $\uparrow$
3DGP [38] 0.47 18.5
IVID [54] 1.23 19.2
G3DR [30] 0.39 36.5
Our F3D-Gaus 0.16 40.4
Table 2: Geometry comparison between our method and three state-of-the-art methods on ImageNet $256^{2}$. Our F3D-Gaus outperforms all three methods in both depth accuracy and NFS.
Method FID $\downarrow$ IS $\uparrow$
IVID [54] 128x 14.1 61.4
G3DR [30] 128x 13.0 136.4
Our F3D-Gaus 128x 1.2 202.8
@@
Table 3: Fidelity comparison with two state-of-the-art methods on ImageNet $128^{2}$. Our F3D-Gaus outperforms both IVID and G3DR at this resolution.
@@
Qualitative results. Fig. 5 showcases the 3DGS renderings from our model in both canonical and novel views across a wide range of categories. These categories include slender-legged insects, animals, household items, detailed objects, and large structures, demonstrating its ability to handle diverse data and produce high-quality results. Fig. 6 provides additional qualitative results on three single-class datasets, further illustrating the effectiveness of our approach. In addition to accurately reconstructing the original view, our method generates plausible and coherent novel views.
@@
Out-of-domain samples. Fig. 8 shows the results of our method on more complex, scene-level inputs that fall outside the scope of the ImageNet dataset. These out-of-domain samples, which include intricate indoor scenes and multi-object environments, introduce a significantly higher complexity compared to the object-centric images in ImageNet. Despite the challenge, our model trained on ImageNet at a resolution of $256^{2}$ successfully produces high-quality novel view renderings, including images, depth maps, and normal maps across different viewpoints, demonstrating its impressive generalization capabilities.
@@ 4.3 Ablation Study
Overall analysis. Fig. 7 illustrates how different components of our method affect the metrics as the rendered view angle increases. For efficiency, the metrics in Fig. 7 are computed on a subset of ImageNet at a resolution of 128. Specifically, the three subplots in Fig. 7 correspond to three metrics: IS, FID, and NFS. Each subplot contains three curves, representing different variants of our framework. The green line represents the baseline, which uses only the pixel-aligned representation without cycle-consistent training or second-stage refinement. In this variant, there is no complementary aggregation, and $\mathcal{L}_{\text{cycle}}$ is removed from Equ. 11, and $\mathcal{L}_{\text{novel}}$ is applied to $\tilde{I_{1}}$ and $\tilde{D_{1}}$, rather than $\hat{I_{1}}$ and $\hat{D_{1}}$. The blue line represents the model with cycle-consistent training, while the red line represents the full model, which includes both cycle-consistent training and second-stage refinement.
Figure 6: Qualitative visualization of rendered images and depth maps on the SDIP Elephants [27], LSUN Horses [61], and Dogs [27] datasets.
@@
Pixel-aligned 3DGS. From the three subplots in Fig. 7, we observe that at a 0-degree angle, the differences between the three variants are minimal, with the baseline (green line) even achieving the highest IS and NFS. This suggests that the generalized 3DGS representation alone can produce high-quality frontal view reconstructions. However, the baseline’s performance degrades significantly as the yaw angle increases. For rendering angles sampled between [-30°, 30°], the baseline’s IS drops to 218.0, FID rises to 11.3, and NFS declines sharply. This indicates that, while the baseline performs well for frontal views, it struggles to represent texture and geometry effectively in novel views.
Figure 9: Demonstration of the effectiveness of cycle-consistent self-supervised training. The comparison between (a) and (b) clearly highlights that our cycle-consistent training strategy significantly mitigates ghosting artifacts.
Cycle-consistent training. In contrast, the blue line, which includes cycle-consistent training, shows more stable performance as the viewing angle increases, with only a slight decline in metrics. This demonstrates that the cycle-consistent training helps our method better leverage monocular datasets to learn 3D perception capabilities.
@@
Fig. 9 illustrates the effect of our cycle-consistent training strategy on the generated novel view results. For two input images, the right side shows the rendered images, depth maps, and normal maps from novel views. Without the cycle-consistent training strategy, the novel views exhibit noticeable ghosting artifacts. For example, for input 1, the novel view of the mug shows a faint outline of an additional edge, and for input 2, the head appears to have multiple overlapping layers. These artifacts suggest that the 3DGS representations derived from different viewpoints are not well-aligned, which contradicts our goal of learning a consistent 3D-aware representation. However, after applying the cycle-consistent training strategy, the rendered results become much more coherent, and the ghosting artifacts are noticeably fixed. This demonstrates that the cycle-consistent strategy helps produce more consistent and aligned 3D representations across different views.
Geometry-guided refinement. In Fig. 7, comparing the blue line and the red line in three subplots, we can find that when we add the second-stage geometry-guided refinement, the red line shows obvious improvements, achieving the best results among all the three metrics. This suggests that the second-stage refinement further enhances the model’s image and geometry rendering. Additionally, the fine-tuning refinement does not affect the multi-view consistency established by the cycle training. The red line shows only minor degradation as the yaw angle increases. This indicates that the second-stage refinement further strengthens the model’s ability to maintain high-quality results, even as the viewing angle becomes more challenging.
@@ 5 Conclusion
In this supplementary material, we provide more additional experiments in Sec. A. We also present a video demo for more qualitative results as shown in Sec. B.
@@ Appendix A Additional Experiments
A.1 Quantitative Fidelity Evaluation
We provide additional quantitative evaluations on three datasets: LSUN Horses [61], SDIP Dogs, and Elephants [27]. The results at a resolution of 256 are summarized in Tab. 4. Our method achieves better FID and NFS scores compared to the previous state-of-the-art approaches, indicating that our method generates more realistic images with smoother depth continuity.
Dogs Horses Elephants
[…]
EG3D [3] 9.83 11.91 2.61 13.34 3.15 2.59
EpiGRAF [37] 17.3 3.53 5.82 9.73 7.25 12.9
IVID [54] 14.7 N/A 10.2 N/A 11.0 N/A
G3DR [30] 8.37 36.89 5.64 36.2 5.30 35.6
Our F3D-Gaus 2.53 41.22 1.14 39.67 2.96 38.92
Table 4: Quantitative evalution on the LSUN Horses [61], SDIP Dogs, and Elephants [27] datasets. Our F3D-Gaus outperforms all the previous methods on both FID and NFS.
A.2 Quantitative Efficiency Evaluation
Method Training Time (A100 days) $\downarrow$ Inference Time (s) $\downarrow$
IVID [54] N/A 20+128
G3DR [30] 14.5 0.9+3.6
Our F3D-Gaus 13.0 0.8+1.1
Table 5: Efficiency comparison between our method and two state-of-the-art methods. Our F3D-Gaus are tested at the resolution of $256^{2}$, while the results of IVID and G3DR are at the resolution of $128^{2}$.
In Tab. 5, we compare the efficiency of our method with IVID and G3DR. Training time is reported in A100 days, while the inference time is measured in seconds. Our approach reduces the training time from 14.5 A100 days (G3DR) to 13 A100 days. The ”inference time” column details the time as ”Time required to obtain the 3D representation + Time for rendering 128 novel views.” Benefiting from the efficiency of our pixel-aligned GS, our method achieves the fastest runtime. While G3DR’s training and inference times are close to ours, their model only supports $128^{2}$ input images. To produce $256^{2}$ outputs, G3DR requires an additional super-resolution step after inference, which introduces extra time not reflected in Tab. 5. In contrast, our method enables direct and efficient training and testing at a native $256^{2}$ resolution, eliminating the need for post-processing.
@@ A.2 Quantitative Efficiency Evaluation
A.3 More Qualitative Results
A.4 Mesh Extraction
Fig. 11 shows the results of mesh extraction from the predicted 3DGS. Our F3D-Gaus renders depth maps as GOF [63] for supervision. Thus, after we get the predicted 3DGS from a single image via F3D-Gaus, we could generate meshes utilizing GOF’s mesh extraction method (tetrahedral grid generation combined with binary search). Fig. 11 demonstrates the capability of our method to predict mesh from a single image via a geometry-driven pipeline leveraging 3DGS. Please note that the mesh is directly derived from the 3DGS representation and does not rely on image-based optimization methods.
A.5 Ablation on Depth Estimation
@@ Appendix B Video Demo
As a supplement to Sec. A.3, we have included a video demo to showcase the quantitative results in a video format. Please check our project page for more details. As shown in Fig.10, we use Real-ESRGAN [49] to upsample $128^{2}$ outputs of G3DR [30] to $256^{2}$.

### v3

@@ 3.3 Geometry-guided Texture Refinement
We found that although our cycle-aggregative self-supervised strategy improves 3D representations in novel views, there are still artifacts near edges with significant viewpoint changes. To address this, we introduce a fixed video in-painting model [66] in the second training stage for artifact localization and additional refinement.
Artifacts localization using normal map. When rendering novel view images from $\hat{GS_{0}}$, we also render the alpha map and calculate the normal map from the depth. We found that alpha and normal maps effectively help identify artifacts. Specifically, if the alpha value is low and the angle between the normal and viewing direction is small, it is likely an artifact. This can be formalized as follows:
@@
where $I_{\text{in}}^{m}$ represents the in-painted image for the $m$-th frame, $\hat{I^{m}}$ is the rendered image, and $M^{m}$ is the mask generated from Equ. 11. Fig. 4 (b) shows one view of the in-painted results. Then, the corresponding video in-painting loss is defined as:
@@ 4.1 Experimental Setup
Dataset. We conducted experiments on ImageNet [9], Dogs [27], SDIP Elephants [27], and LSUN Horses [58], following previous methods. ImageNet contains 1000 categories, while the others are single-class images. For fair comparison, we use pseudo-ground-truth depth from LeReS [56], though our approach supports other depth estimators. Following G3DR [30], we train on a filtered subset (420K images) and evaluate on the full set (1.2M images). Most of our experiments are conducted on ImageNet.
Metrics. For the generated images, we compute Fréchet Inception Distance (FID) [13] and Inception Score (IS) on the full ImageNet training set. For the predicted depth maps, we compute the Non-Flatness Score (NFS) [35], a no-reference depth quality metric that evaluates the continuity by analyzing the histogram distribution of the depth map.
[…]
Settings. We evaluate our model with two different input types and compare it against previous methods. First, we use ImageNet images as input to reconstruct the original view for comparison with G3DR. Second, we used 10k images sampled by the IVID [52] to render novel views, to verify the effectiveness of the entire two-stage pipeline with the same 2D conditional image generation module. Yaw and pitch angles of novel camera are sampled from Gaussian distributions with variances of 0.3 and 0.15, respectively.
@@ 4.2 Main Results
Quantitative results. Fig. 5, Tab. 1, and Tab. 2 present our qualitative and quantitative results on ImageNet. In Tab. 1, we compare G3DR’s “single-view 3D-aware model” $f_{\textit{trigen}}$ with ours, using the full ImageNet as input. The outputs are then upscaled to $256^{2}$ via Real-ESRGAN [46] for comparison. The top part of Tab. 1 shows our method significantly outperforms G3DR in image fidelity and depth consistency. To eliminate the potential gap introduced by Real-ESRGAN, we also compare our $128^{2}$ model with $f_{\textit{trigen}}$ using metrics directly at 128 resolution. Tab. 1 confirms that our method outperforms G3DR at both resolutions.
We further adopt another setting to evaluate the quality of generated novel views images by “2D conditional generation + 3D generation”. In Tab. 2, we use 10K sampled images from IVID-stage1 with class labels as input. Tab. 2 shows that our method achieves the best FID/NFS and competitive IS, highlighting the advantage of pixel-aligned GS in 3D-aware generation.
We provide additional quantitative evaluations on LSUN Horses [58], SDIP Dogs, and Elephants [27], summarized in Tab. 4 at 256 resolution. Our method outperforms G3DR [30] with better FID and NFS, indicating that our method generates more realistic images with smoother depth continuity. Since G3DR’s pre-trained models for these datasets are unavailable, we train the $f_{\textit{trigen}}$ models using the released code, and upsample G3DR [30]’s $128^{2}$ outputs to $256^{2}$ with Real-ESRGAN [46].
Method FID $\downarrow$ IS $\uparrow$ NFS $\uparrow$
G3DR-$f_{\textit{trigen}}$ + RealESRGAN ($256^{2}$) 24.9 78.2 34.1
Our F3D-Gaus ($256^{2}$) 1.6 308.6 40.4
G3DR-$f_{\textit{trigen}}$ ($128^{2}$) 14.5 86.3 34.0
Our F3D-Gaus ($128^{2}$) 1.2 202.8 40.6
@@
Table 1: Comparison on ImageNet with G3DR. Our F3D-Gaus surpasses the state-of-the-art 3D generation methods G3DR [30] in FID, IS, and NFS metrics in both $128^{2}$ and $256^{2}$ resolution.
Method FID $\downarrow$ IS $\uparrow$ NFS $\uparrow$
IVID [52] 128x 16.5 216.4 25.7
G3DR [30] 128x 15.8 135.3 36.1
Our F3D-Gaus 128x 15.5 182.7 37.4
Table 2: Comparison with state-of-the-art methods on 10k sampled images. F3D-Gaus outperforms in two of three key metrics, highlighting its strengths in fidelity and geometry consistency.
Method Training $\downarrow$ Inference (s) $\downarrow$
IVID [52] N/A 20+128
G3DR [30] 14.5 0.9+3.6
Our F3D-Gaus 13.0 0.8+1.1
Table 3: Efficiency comparison with state-of-the-art methods. Our F3D-Gaus achieves the fastest training (measured in A100 days) and inference speed.
@@
Qualitative results. Fig. 5 showcases 3DGS renderings from our model in both canonical and novel views across a wide range of categories, including slender-legged insects, animals, household items, detailed objects, and large structures. These results highlight our model’s ability to handle diverse data and generate high-quality outputs. Fig. 6 provides additional qualitative results on three single-class datasets, further illustrating the effectiveness of our approach. Beyond accurately reconstructing the original view, our method generates plausible and coherent novel views.
@@
Out-of-domain samples. Fig. 8 shows the results of our method on more complex, scene-level inputs that fall outside the scope of ImageNet. The out-of-domain samples, including intricate indoor scenes and multi-object environments, introduce significantly higher complexity than ImageNet’s object-centric images. Despite the challenge, our model trained on ImageNet at $256^{2}$ resolution successfully produces high-quality novel view renderings, including images, depth, and normal maps across various viewpoints, demonstrating its impressive generalization capabilities.
Efficiency. Tab. 3 compares our method’s efficiency with IVID and G3DR. Our approach reduces the training time from 14.5 to 13 A100 days. Inference time is reported as ”Time to obtain the 3D representation + Time to render 128 novel views.” Leveraging pixel-aligned GS, our method achieves the fastest runtime. While G3DR’s training and inference times are close to ours, it only supports $128^{2}$ inputs, which requires an additional super-resolution step to upscale $128^{2}$ outputs to $256^{2}$. Our pixel-aligned Gaussian design enables direct training and testing at $256^{2}$, eliminating post-processing overhead not reflected in Tab. 3.
Method ($256^{2}$) Dogs Horses Elephants
[…]
G3DR-$f_{\textit{trigen}}$ + RealESRGAN 14.11 37.16 11.65 32.31 24.39 33.02
Our F3D-Gaus 2.53 41.22 1.14 39.67 2.96 38.92
Table 4: Quantitative evaluation on the LSUN Horses [58], SDIP Dogs, and Elephants [27] datasets.
Figure 6: Qualitative results of rendered images and depth maps on SDIP Elephants [27], LSUN Horses [58], and Dogs [27].
@@ 4.3 Ablation Study
Overall analysis. Fig. 7 shows how different components of our method affect various metrics as the view angle increases. For efficiency, these metrics are computed on a ImageNet subset at 128 resolution. The three subplots in Fig. 7 correspond to IS, FID, and NFS, each with three curves representing different variants. The green line represents the baseline, using only the pixel-aligned representation, without cycle-aggregative training or second-stage refinement. In this variant, no complementary aggregation is applied, $\mathcal{L}_{\text{cycle}}$ is removed from Equ. 10, and $\mathcal{L}_{\text{novel}}$ is applied to $\tilde{I_{1}}$ and $\tilde{D_{1}}$, rather than $\hat{I_{1}}$ and $\hat{D_{1}}$. The blue line represents the model with cycle-aggregative training, while the red line represents the full model, incorporating both cycle-aggregative training and second-stage refinement.
@@
Pixel-aligned 3DGS. Fig. 7 shows that at 0° angle, all three variants perform similarly, with the baseline (green line) even achieving the highest IS and NFS. This suggests that the generalized 3DGS representation alone can produce high-quality frontal view reconstructions. However, as the yaw angle increases, the baseline’s performance degrades sharply. For angles sampled from [-30°, 30°], its IS drops to 217.8, FID rises to 11.7, indicating that it struggles to preserve texture and geometry effectively in novel views.
@@
Cycle-aggregative training. In contrast, the blue line, which incorporates cycle-aggregative training, performs better at wider viewing angles. For instance, at [-30°, 30°] yaw angle, its FID drops from 11.7 to 8.0 (Fig. 7 (b)), showing that cycle-aggregative training improves 3D perception using monocular datasets. Fig. 9 further illustrates its effect on novel views. The right side shows rendered images, depth, and normal maps for two input images. Without cycle training, novel views exhibit noticeable ghosting artifacts, e.g., input 1’s mug has a faint extra edge, and input 2’s head appears as multiple overlapping layers. These artifacts indicate that 3DGS representations derived from different viewpoints are not well-aligned, which contradicts our goal of learning a consistent 3D-aware representation. However, after applying cycle-aggregative training, the rendered results become much more coherent, eliminating ghosting artifacts. This demonstrates its effectiveness in producing consistent and aligned 3D representations across views.
Figure 9: Demonstration of the effectiveness of cycle-aggregative self-supervised training. The comparison between (a) and (b) clearly highlights that our cycle-aggregative training strategy significantly mitigates ghosting artifacts.
Geometry-guided refinement. In Fig. 7, comparing the blue and red lines, we observe that the second-stage geometry-guided refinement obviously improves performance. The red line achieves the best results among all three metrics, indicating that this refinement further enhances both image and geometry rendering.
@@ 5 Conclusion
In this supplementary material, we provide more additional experiments In this supplementary material, we provide more additional experiments in Sec. A. We also present a video demo for more qualitative results as shown in Sec. B.
@@ Appendix A Additional Experiments
A.1 More Qualitative Results
@@
A.2 Mesh Extraction
Fig. 11 shows the results of mesh extraction from the predicted 3DGS. Our 3DGS representation is based on GOF [60] for accurate depth and normal rendering, which naturally supports surface reconstruction. After we get the predicted 3DGS from a single image, we extract meshes using GOF’s tetrahedral grid generation combined with a binary search. Fig. 11 demonstrates the capability of our method to predict mesh from a single image via a geometry-driven pipeline leveraging 3DGS. Please note that the mesh is directly derived from the 3DGS representation and does not rely on image-based optimization methods.
A.3 Ablation on Depth Estimation
@@
A.4 Why no conditional generation model?
The pipeline of state-of-the-art methods G3DR [30] and IVID [52] can be divided into a “class-conditioned RGBD generation model” and a “3D-aware generation model with RGBD input”, where we focus on the latter. Actually, training an additional conditional generation model to fit the ImageNet distribution is straightforward with existing frameworks like GANs, Diffusion models, and VARs. Including a preceding RGBD generation (like G3DR and IVID) in evaluation introduces confounding factors that blend both models’ performance, detracting from our core contribution. Although we did not train a conditional generation model, our method still shows extensibility to images generated by existing models within the ImageNet distribution. As shown in Fig. 14, our method performs well on in-domain generated images (from VAR [41] trained on ImageNet) and even out-of-domain images (FFHQ [20]), demonstrating its generalization ability.
Figure 14: Qualitative results on VAR [41] output (left) and FFHQ images (right).
@@ Appendix B Video Demo
As a supplement to Sec. A.1, we have included a video demo to showcase the quantitative results in a video format. As shown in Fig.10, we use Real-ESRGAN [46] to upsample $128^{2}$ outputs of G3DR [30] to $256^{2}$.